OpenStudy (anonymous):

A ball contains 40 balls numbered 1 through 40. If 1 through 10 are considered lucky and 2 balls are drawn from the bowl without replacement, find the probability that neither balls are lucky. Can someone help me answer this ?????

OpenStudy (anonymous):

"A ball contains 40 balls" - what kind of ball is it?

OpenStudy (anonymous):

probability first ball is "not lucky" is \(\frac{30}{40}=\frac{3}{4}\) probability second ball is not lucky "given" that the first ball is not lucky is \(\frac{29}{39}\) reasoning that there are 39 balls left left and of those 29 are not lucky probability both things occur is the product of these two, namely \[\frac{3}{4}\times \frac{29}{39}\] it is an urny ball

OpenStudy (anonymous):

Do you mean bowl?

OpenStudy (anonymous):

a bowl looking to get lucky

OpenStudy (anonymous):

yes i meant bowl